OpenStudy (itsmichelle29):

help please medal and fan Solve the equation on the[0,2π). interval ​[0,2piπ​). 18sinx=8cos^2 x−15

hartnn (hartnn):

hint : if you write cos^2 x in terms of sin^2 x, you will get a quadratic equation in sin x :)

OpenStudy (itsmichelle29):

I dont understand

hartnn (hartnn):

how will you write cos^2 x in terms of sin^2 x? any pythogorean identity?

OpenStudy (itsmichelle29):

i am completely lost .. i really need that step to step process

hartnn (hartnn):

did you know about \(\sin^2 x+\cos^2 x =1\) ?

OpenStudy (itsmichelle29):

Yes

OpenStudy (itsmichelle29):

1-sin2x=cos2x

hartnn (hartnn):

what is cos^2 x from that ?

OpenStudy (itsmichelle29):

theres that one too i think

hartnn (hartnn):

you mean \(\cos^2 x = 1- \sin^2 x\) ? yes, you need to use this and replace cos^2 x by the right side

OpenStudy (itsmichelle29):

yes thats what i meant :)

hartnn (hartnn):

18sinx=8(1-sin^2 x)−15 let u = sin x

OpenStudy (itsmichelle29):

18u=8(1-u^2)-15

OpenStudy (itsmichelle29):

this is really stressing

OpenStudy (itsmichelle29):

i dont understand

hartnn (hartnn):

but did you get this: 18sinx=8(1-sin^2 x)−15 ?

OpenStudy (itsmichelle29):

yes u changed it it ... and it equals the same

hartnn (hartnn):

now i just made a substitution, for simplicity writing sin x as u

hartnn (hartnn):

18sinx=8(1-sin^2 x)−15 18u=8(1-u^2)−15 this is a quadratic in u

OpenStudy (itsmichelle29):

Okay

hartnn (hartnn):

can you solve quadratics?

OpenStudy (itsmichelle29):

no

OpenStudy (itsmichelle29):

okay